Need help did this problem and got it wrong:
Factor the expression.
x(x - y - z) + y(x - y - z)
put answer x to the second -2x-y to the second -yz

Answer by Theo(13342)

x * (x - y - z) = x^2 - xy - xz (first part)

y * (x - y - z) = yx - y^2 - yz (second part)

since yx is the same as xy, this second part becomes:

y * (x - y - z) = xy - y^2 - yz (second part)

add first part and second part together and you get:

(x^2 - xy - xz) + (xy - y^2 - yz) which becomes:

x^2 - xy - xz + xy - y^2 - yz.

The + xy and the - xy cancel out so the equation becomes:

x^2 - xz - y^2 - yz

your answer was:

x^2 - 2x - y^2 - yz

not sure where you got the 2x from.

I believe that should have been zx or xz.